OpenStudy (precal):

Is there a trick to finding limits as x approaches infinity using vertical asymptotes?

OpenStudy (precal):

I am given a rational function. I know how to find the vertical asymptote but the question is asking about limits

OpenStudy (precal):

\[f(x)=\frac{ 2x^2+x-3 }{ x^2+4x-5 }\]

OpenStudy (precal):

the function above has a vertical asymptote at x=-5 because

OpenStudy (precal):

limit as x approaches -5 from the left side is equal to infinity

OpenStudy (precal):

I can't use a graphing calculator, because if I could then it would be easy to see infinity

OpenStudy (precal):

so can this be done without a graphing calculator

zepdrix (zepdrix):

Yah there's a trick :) Infinite limits tells us about `horizontal asymptotes`. If the `largest` degree of top and bottom are `equal`, then as x->infty, the limit is approaching the ratio of the leading coefficients. So in your example, we would have a horizontal asymptote at 2/1.

zepdrix (zepdrix):

The other two cases are when, the denominator is `greater` degree than the numerator, in which case the asymptote is y=0 (the x-axis). Think of 1/x (larger degree in the bottom) asymptotes are both axes, yes?

zepdrix (zepdrix):

And the last case is when the top is larger, I think we get no asymptote if I remember correctly....

OpenStudy (precal):

I know those tricks using the horizontal asymptotes. I have to use vertical asymptotes as a justification

zepdrix (zepdrix):

Ohhh lol I musta misread the question haha sorry >.< So we simply factor the denominator, yes?

OpenStudy (precal):

yes I factored both top and bottom, one factor cancel out and I am left with the vertical asymptote

OpenStudy (precal):

this is an ap calculus question, there is always some twist to how they ask and answer questions

OpenStudy (amistre64):

x approach infinity is a horizontal/oblique asymptote ... has nothing to do with vertical asymptotes

OpenStudy (precal):

yes, I know. Maybe I should write the question again with all of its choices.

OpenStudy (amistre64):

i cant think of any rational reason why knowing what happens at x=-5 would tell us anything about what happens as x moves towards infinity.

OpenStudy (precal):

The function \[f(x)=\frac{ 2x^2+x-3 }{ x^2+4x-5 }\] has a vertical asymptote at x=-5 because

OpenStudy (amistre64):

becuase it has no controling factor at x=-5

zepdrix (zepdrix):

I'm not exactly sure what you're asking I guess >.< But lemme try to give some info that could maybe be useful. \[\Large\rm \lim_{x\to-5}\frac{ 2x^2+x-3 }{ x^2+4x-5}=\lim_{x\to-5}\frac{(x-1)(2x+3) }{(x-1)(x+5)}\]And then I guess you would have to split this into uhhh one-sided limits or whatever those are called :P\[\Large\rm \lim_{x\to-5^+}\frac{(x-1)(2x+3) }{(x-1)(x+5)}\]So from the positive side... that means the value is slightly larger than -5 (like -4.9999), so we'll end up with:\[\Large\rm \lim_{x\to-5^+}\frac{(x-1)(2x+3) }{(x-1)(x+5)}=\frac{-7}{0^+}\]The denominator is slightly larger than 0. So overall we're approaching `negative infinity` because of our signs.

OpenStudy (amistre64):

IF there was a control factor: (x+5)/(x+5), then the limit as x approaches -5 would even out to 1 and youd have a hole, a removable discontinuity.

zepdrix (zepdrix):

And we end up approaching `positive infinity` from the left side because we get a little less than 0 in the denominator from that side. Oh boy messages are lagging :P

OpenStudy (precal):

I have 5 choices A) limit as x approaches -5 from the right of f(x) = infinity B) limit as x approaches -5 from the left of f(x) = negative infinity C) limit as x approaches -5 from the left of f(x) = infinity D) limit as x approaches negative infinity of f(x)=-5 E) f(x) does not have a vertical asymptote at x=-5

OpenStudy (precal):

Solution is C just trying to understand why this is so without the use of a graphing calculator

OpenStudy (precal):

thanks for you help. Some of these questions do allow the use of a graphing calculator while others do not.

OpenStudy (precal):

@zepdrix you were on the right track I found the explanation buried in other pages they did take numbers to the left or right side and compare the signs example positive over negative to determine negative infinity or positive over positive for positive infinity Thanks for your help